# Logic Augmented Generation

### v1

@@ 5 Conclusions and future work
In this work we introduce Logic Augmented Generation (LAG), a novel solution that integrates Semantic Knowledge Graphs (SKGs) with Reactive Continuous Knowledge Graphs (RCKGs) to enforce logical consistency and blend symbolic and continuous reasoning. RCKGs are framed as a neuro-symbolic approach that uses Large Language Models (LLMs) to extract and represent knowledge, incorporating tacit insights and contextual nuances. This hybrid model enables context-aware knowledge generation and addresses challenges like formalising tacit knowledge, enhancing prompt engineering, and aligning plausibility- and truth-preserving logic. The potentiality of LAG is illustrated by using collective intelligence as a case study. Future work will study SKG-RCKG integration, advance prompt strategies, and expand operators for representing tacit semantics.

### v2

@@ 5 Conclusions and future work
In this work we introduce Logic Augmented Generation (LAG), a novel solution that integrates Semantic Knowledge Graphs (SKGs) with Reactive Continuous Knowledge Graphs (RCKGs) to enforce logical consistency and blend symbolic and continuous reasoning. RCKGs are products of a neuro-symbolic method that uses Large Language Models (LLMs) to extract and represent knowledge, incorporating tacit insights and contextual nuances. This hybrid model enables context-aware knowledge generation and addresses challenges like formalising tacit knowledge, enhancing prompt engineering, and aligning plausibility- and truth-preserving logic. The potentiality of LAG is illustrated by using collective intelligence as a case study. Future work will study SKG-RCKG integration, advance prompt strategies, and expand operators for representing tacit semantics.